Hello,
I'm having difficulties setting up this word problem:
Sam bought 50 stamps at 14 and 22 cents each. He paid the postal clerk
$10.04. How many 22-cent stamps did he buy?
I tried to set the problem up like this:

x+y=50 letting the x=14 and y=22
y=x-50
14x + 22y = 50
14x + 22(50-x)=50
14x + 1100 -22x = 50
-8x = -1050
-8x/-8 = -1050/8
x= 131.25
Not sure where I went wrong. I appreciate any help. Thank you.

Answer by rwm(914)

x+y=50
let x=14cent stamps and y=22 cent stamps
x+y=50
14x + 22y = 1004
try again